OpenStudy (anonymous):

write the slope-intercept equation for the line that passes through (10, -4) and has a slope -3/10

OpenStudy (anonymous):

you need to calculate the y-intercept by plugging all the given values in y=mx + b -4 = (-3/10)(10) +b -4 = -3 + b -1 = b :- y = (-3/10)x - 1 alternative: 10y = -3x - 10